OpenStudy (anonymous):

find second derivative if (x-5)^5+(y-2)^5=64

OpenStudy (amistre64):

f1{x,y} = 5(x-5)^4 f11{x,y} = 20(x-5)^3

OpenStudy (amistre64):

f2{x,y} = 5(y-2)^4 f22{x,y} = 20(y-2)^3

OpenStudy (anonymous):

5(x-5)^4 + 5(y-2)^4 dy/dx =0 ist derivative

OpenStudy (amistre64):

right, so long as we imply that y is a function of x

OpenStudy (anonymous):

yes,, maybe we can let snenlaha do the 2nd derivative hehehehe

OpenStudy (amistre64):

maybe :)

OpenStudy (anonymous):

i couldn't continue from the first to the 2nd derivative

OpenStudy (anonymous):

but do you know the process of the ist that i did?

OpenStudy (anonymous):

yes i do and i got it the same as yours

OpenStudy (anonymous):

ok try the second so that you cn have confidence in yourself

OpenStudy (anonymous):

do i need to make dy/dx a subject of the formula

OpenStudy (anonymous):

try to arrange them first, start with dy/dx =

OpenStudy (amistre64):

you can; y' just derives to y''; just be sure to use it as such in its product rule

OpenStudy (anonymous):

yes thats the other way doing it,, either one will arrive at the right solution

OpenStudy (amistre64):

to implicit it again would be easier; no quotient ruls to fight with i believe

OpenStudy (anonymous):

ok thats rigth....try implicitly ..

OpenStudy (amistre64):

5(x-5)^4 + 5(y-2)^4 y' =0 20(x-5)^3 + 5(y-2)^4 y'' + 20(y-2)^3 y' y' = 0 then if it works out that way :)

OpenStudy (anonymous):

Dx(5(x-5)^4) + Dx(5(y-2)^4 dy/dx) =Dx(0).......ok amistre did it now hehehehe

OpenStudy (amistre64):

i got no idea if its right :)... other than the obvious typos that is -20[(x-4)^3 + (y-2)^3 y'^2] y'' = ------------------------- right? 5(y-2)^3

OpenStudy (amistre64):

-20/5 reduces..... and y' = whatever it equals lol

OpenStudy (anonymous):

the answer is suppose to be: y''=(-256(X-5)^3)/(y-2)^9

OpenStudy (anonymous):

dy/dx=-(x-5)^{4}/(y-2)^{4}\[d ^{2}y/dx ^{2}=-256(x-5)^{3}/(y-2)^{5}\]

OpenStudy (amistre64):

it prolly simplifies to that then lol...

OpenStudy (anonymous):

yeah ,... thats the answer to it Snenhlala....try to practice till you arrive to that answer

OpenStudy (anonymous):

sorry the denominator for the second derivative should read (y-2)^9

OpenStudy (anonymous):

did you used product rule or quotient rule to find 2nd derivative

OpenStudy (anonymous):

i did mine in quotient rule

OpenStudy (anonymous):

use the quotient rule and sub for dy/dx

OpenStudy (anonymous):

mulitply each term by (y-2)^4 then factor -4(x-5)^3(y-2)^3 this will leave (x-5)^5+(y-2)^4 which equals 64 (from the original equation)

OpenStudy (anonymous):

should read (x-5)^5+ (y-2)^5 which equals 64

OpenStudy (anonymous):

should i start to multiply from the 1st derivative or from the origin equation?

OpenStudy (anonymous):

multiply the second derivative

OpenStudy (anonymous):

by (y-2)^4

OpenStudy (anonymous):

y'= -(x-5)^4 / (y-2)^4 -4(y-2)^4 (x-5)^3 - 4(x-5)^4 (y-2)^3 y' y"= ----------------------------------- now subst. y ' in this equqtion ((y-2)^4)^2 -4(y-2)^4 (x-5)^3 - 4(x-5)^4 (y-2)^3 [-(x5^4)/(y-2)^4] y"= ------------------------------------------------------------- ((y-2)^4)^2 -4(y-2)^4 (x-5)^3 + 4(x-5)^8 (y-2) y"= ----------------------------------- (y-2)^8

OpenStudy (anonymous):

-4(y-2)^4 (x-5)^3 + 4(x-5)^8/(y-2) y"= ----------------------------------- ... sorry its divided by y-2 (y-2)^8

OpenStudy (anonymous):

-4(y-2)^5 (x-5)^3 + 4(x-5)^8 y"= ----------------------------------- (y-2)^9

OpenStudy (anonymous):

-4 (x-5)^3 [(y-2)^5 + (x-5)^5] y"= ----------------------------------- (y-2)^9

OpenStudy (anonymous):

-4 (x-5)^3 [64] y"= -------------- (y-2)^9 57 seconds ago

OpenStudy (anonymous):

-256 (x-5)^3 y"= -------------- (y-2)^9

OpenStudy (anonymous):

well i think thats it hehehe

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

w.c.

OpenStudy (anonymous):

did you arrived in the same procedure?

OpenStudy (anonymous):

and of course the same answer? lol

OpenStudy (anonymous):

yes i did arrive but it too tricky

OpenStudy (anonymous):

yeah kind of hehehe,,